5x-6=4x-2 please please please tell me I can’t understabd this please

Grade:7

1 Answers

Vijay
2 years ago
Here , Given equation :
  • It means that  what is the value of x to satisty this equation.
5x-6=4x-2 
our next step is to bring all expression in one side -
so , 5x-6-4x+2=0
we get , x-4=0 after that , x=4
  • so x=4 , is the correct answer for this expression.